# arXiv:2603.00283v1[eess.SY] 27 Feb 2026

### v1

@@ Abstract
thanks: This work is supported by AFOSR through the grant FA9550-21-1-0411, NASA through the grant 80NSSC20M0229 and NSF through the AI Institute: Planning grant #2020289. The material is this paper was partially presented at the 2024 American Control Conference (ACC), 10-12 July 2024, Toronto, ON, Canada. Corresponding author Pan Zhao.
@@ 2 Problem Statement
It is possible to extend the results to allow $w(t)$ to be state-dependent as well, $w(t,x)$, provided the boundedness assumption (3) holds. To simplify the exposition, we limit the derivations in this paper to time-dependence only.
@@ 3 Overview of the UC-MPC Framework
$\!\dot{x}(t)\!$ $=\!{A_{m}}x(t)\!+\!Bu_{\text{opt}}(t)\!+\!B(u_{\textup{a}}(t)\!+\!f(t,\!x(t)\!)\!)\!+\!B_{u}w(t)$
$y(t)$ $=Cx(t),\quad x(0)=x_{0}$ (8)
@@ 3.1 Equivalent Nominal System and Connection with Tube MPC
Why use two nominal systems? We first introduce the nominal system in its original form (9) because the performance analysis of $\mathcal{L}_{1}$}AC—specifically the derivation of its uniform performance bounds —relies on comparing this nominal system with the actual closed-loop dynamics, where their difference arises solely from system uncertainties. These performance bounds will be discussed in detail in Section 5. The transformation in (12) reveals a close connection between tube MPC and our proposed UC-MPC. In both frameworks, the feedback term $K_{x}(x-x_{\textup{n}})$ serves to keep the actual trajectory within a bounded neighborhood of the nominal one, while the adaptive component $u_{a}$ in UC-MPC actively compensates for matched uncertainties to further improve the performance. Owing to this adaptive compensation, the resulting deviation tube in UC-MPC is expected to be significantly tighter than that of the standard tube MPC. This reformulation also allows the MPC input $u_{\text{opt}}$ in (9) to be computed directly from (12), which defines a standard uncertainty-free MPC problem same as [22, 4, 6]. Consequently, the stability and recursive feasibility of the nominal MPC can be established using existing MPC analysis tools, which will be discussed in detail in Section 6.2. To summarize, the nominal system in (9) and control law in (7) serve as the foundation for all performance bounds analysis in Section 5, while the equivalent transformed form in (12) and control law in (13) are introduced for formulating the nominal MPC problem and highlighting the connection between UC-MPC and tube MPC [16], as well as for the stability and recursive feasibility analysis of the nominal MPC problem within the UC-MPC framework.
@@ 4 Unmatched Uncertainty Attenuation
Consider the following system,
@@
where $x(t)\in\mathbb{R}^{n}$, and $u(t)\in\mathbb{R}^{m}$ represent the state and input vectors, respectively, $w(t)\in\mathbb{R}^{p}$ represents the unmatched uncertainty, $z(t)$ denotes output channels whose bounds we aim to minimize, and matrices $A,B$, and $B_{u}$ are from (1). Note that when the matched uncertainty is perfectly eliminated, the original system (1) is identical to the system in (14a). With the feedback control law $u(t)=K_{x}x(t)$, we achieve the following closed-loop system,
@@
$\begin{bmatrix}(A\!+\!BK_{x})^{\top}P\!+\!P(A\!+\!BK_{x})\!+\!\lambda P&PB_{u}\\ B^{\top}_{u}P&-\mu I\end{bmatrix}<0$ (16)
@@
Multiplying (18) by $\begin{bmatrix}\mathcal{V}^{-1}&0\\ 0&I\end{bmatrix}$ and its transpose from left and right gives (16), where $P=\mathcal{V}^{-1}$ and $K=\mathcal{Y}\mathcal{V}^{-1}$. Similarly, multiplying (19) by $\textup{diag}(\mathcal{V}^{-1},I,I)$ and its transpose from left and right gives (17). Thus, we can leverage Lemma 1 to achieve the internal stability and PPG bound $\beta$. ∎
Based on Lemma 2, we can derive matrices $\mathcal{Y}$ and $\mathcal{V}$ to minimize $\beta$ for a given $\lambda$ by solving the following LMI:
@@
By setting $K_{x}=\mathcal{Y}\mathcal{V}^{-1}$, we achieve the minimized PPG bound $\beta$ for system $H$ and less tightened constraint tightening for channels in $z(t)$. Here, we have to pick a $\lambda$ first since 16 and 17 are only linear if we fix $\lambda$ [25]. For the selection of $\lambda$, a linear search between 0 to some predefined maximum value can be implemented for an improved performance.
@@
Lemma 2 provides one method to synthesize a state-feedback controller $u=K_{x}x$ to minimize the effect of bounded disturbances. There are other methods available in the literature, e.g., [1] and [9], both of which leverage inescapable ellipsoids to over-approximate the reachable sets of the disturbed system 14 and provide necessary and sufficient conditions for ensuring systems stay in prescribed ellipsoids. These methods can be seamlessly integrated into our proposed UC-MPC framework to replace the approach stated in Lemma 2, and can potentially lead to less conservative performance.
@@ 5 $\mathcal{L}_{1}$}AC with Uniform Performance Bounds
Given the uncertain system 8 subject to Assumption 1, the reference system 30 subject to the conditions 22b and 22a with a constant $\gamma_{1}>0$, and an $\mathcal{L}_{1}$}AC defined via 24, 25 and 26 subject to the sample time constraint 29, we have
@@
Given the reference system 30 and the nominal system 9 subject to Assumption 1 and the condition 22a, we have
@@ 5.1 Individual Performance Bounds
Consider the reference system 30 subject to Assumption 1, the nominal system 9, the transformed reference system 35 and transformed nominal system 39 obtained by applying 33 with any $\mathrm{T}_{x}^{i}$ satisfying 34. Suppose that 22a holds with some constants $\rho_{r}$ and $\left\lVert u_{\text{opt}}\right\rVert_{\mathcal{L}_{\infty}}$. Then, there exists a constant $\check{\rho}_{r}^{i}\leq\rho_{r}$ such that 41 holds with the same $\left\lVert u_{\text{opt}}\right\rVert_{\mathcal{L}_{\infty}}$. Furthermore, $\forall t\geq 0$,
@@
Consider the uncertain system 8 subject to Assumption 1, the nominal system 9, and the $\mathcal{L}_{1}$}AC defined via 24, 25 and 26 subject to the conditions 22b and 22a with constants $\rho_{r}$ and $\gamma_{1}>0$ and the sample time constraint 29. Suppose that for each $i\in\mathbb{Z}_{1}^{n}$, 41 holds with a constant $\check{\rho}_{r}^{i}$ for the transformed reference system 35 obtained by applying 33. Then, $\forall t\geq 0$, we have
@@
Lemma 6 introduces a technique to individually bound $x_{i}(t)-x_{\textup{n},i}(t)$ for each $i\in\mathbb{Z}_{1}^{n}$ and $u_{j}(t)-u_{\textup{n},j}(t)$ for each $j\in\mathbb{Z}_{1}^{m}$ through coordinate transformations. Furthermore, in the absence of unmatched uncertainty, by reducing $T$ and increasing the bandwidth of the filter ${\mathcal{C}}(s)$, the value of $\tilde{\rho}^{i}$ ($i\in\mathbb{Z}_{1}^{n}$) can be made arbitrarily small. This implies that, the adaptive system states can be made arbitrarily close to the nominal system states. Additionally, the bounds on $u_{\textup{a},j}(t)$ and $u_{j}(t)-u_{\textup{n},j}(t)$ can be brought arbitrarily close to the bounds on the true matched uncertainty $f_{j}(t,x)$ for $x\in\mathcal{X}_{a}$ and each $j\in\mathbb{Z}_{1}^{m}$. We notice that the low-pass filter in $\mathcal{L}_{1}$}AC decides the tradeoff between performance and robustness. While reducing $T$ can lead to high-frequencies in the estimation loop, an appropriately selected bandwidth and order for the low-pass filter should prevent these high-frequencies to enter the system. Additional details on the role and systematic design of the filter can be found in [7].
@@ 5.2 Scaling for Improved Bounds
where $\bar{B}_{u}=B_{u}\Lambda$ and $\bar{w}$ = $\Lambda^{-1}w$ for some positive diagonal matrix $\Lambda$ s.t. $\left\lVert\bar{w}_{i}(t)\right\rVert_{\mathcal{L}_{\infty}}=b_{w}\;\forall i$. It follows that $B_{u}w=\bar{B}_{u}\bar{w}$, and thus system (47) is identical to the original system (1), and the assumption in (4c) still applies to $\bar{w}(t)$. We can easily achieve the same results in Section 5.1 for the transformed system (47) but with smaller performance bounds. This is because during the calculation of the performance bounds in $\mathcal{L}_{1}$}AC, we only leverage $\mathcal{L}_{\infty}$ bound $b_{w}$. The additional information of individual bounds $b_{w_{j}}$ can help us design $\Lambda$, which makes the elements in $\bar{B}_{u}$ smaller than $B_{u}$ as $b_{w_{j}}\leq b_{w}$, and thus leading to a smaller performance bounds for the transformed system (47) compared with the original system (1). In Section 7.1.5, we will show that with the transformation of $\bar{B}_{u}$ and $\bar{w}$ in (47), which we denote as scaling of $w(t)$, we achieve a smaller performance bound for every state and control input channel provided by $\mathcal{L}_{1}$}AC compared with the original system (1).
@@ 6 UC-MPC: Robust Adaptive MPC via Uncertainty Compensation
2:Find an upperbound range of $u_{\text{opt}}$ and calculate $\left\lVert u_{\text{opt}}\right\rVert_{\mathcal{L}_{\infty}}$ based on 9 with constraints $x_{\textup{n}}\in\mathcal{X}$ and $u_{\textup{n}}\in{\mathcal{U}}$
@@
  • Design a low-pass filter in (21) that satisfies (22).
  • Apply the transformation 33 with $\mathrm{T}_{x}^{i}$ satisfying 34 to achieve $\check{\rho}_{r}^{i}$ that satisfies 41.
@@
The details are summarized in Algorithm 1. It is important to note that, in step 3 of Algorithm 1, we also impose the constraint that both $x_{\textup{r}}(t)$ and $x(t)$ remain within $\mathcal{X}$ for all $t\geq 0$. These constraints can help reduce the size of the uncertainty that needs to be addressed and substantially decrease the conservatism of the proposed approach. Finally, we formulate an MPC problem for the nominal system 12 using tightened constraint sets.
@@ 6.1 Nominal MPC Formulation
where $T_{f}$ is the time horizon, $\l _{f}$ denotes the custom terminal cost function, $l$ denotes the running cost function, and $\mathcal{X}_{f}$ is the custom terminal constraint set. Finally, we can apply the control $\bar{u}_{\text{opt}}$, defined as $\bar{u}_{\text{opt}}(\tau+t)=u^{*}(t)$, to the system, for $0\leq t\leq t_{\delta}$, where $u^{*}(\cdot)$ is the solution to the optimization problem 49 given $x_{\textup{n}}(\tau)$, and $t_{\delta}$ is a sufficiently small sampling time. It is worth noting that $\bar{u}_{\text{opt}}$ is derived solely from the nominal closed-loop dynamics, without relying on the actual system state affected by uncertainties. Consequently, $\bar{u}_{\text{opt}}$ can be precomputed for all time steps offline, prior to the online execution of UC-MPC.
@@ 6.4 Formal Guarantees of UC-MPC
Consider an uncertain system (1) subject to Assumption 1 with the state and input constraints in (6). Consider the control law (13), where $u_{a}$ is the control input of an adaptive controller designed following Algorithm 1, and $\bar{u}_{\textup{opt}}$ is obtained from solving the nominal MPC problem 49. If Assumptions 2 and 3 hold, we have:
@@ 6.5 Computational Complexity Analysis
Compared with the vanilla MPC and tube MPC formulations in [16, 10], the proposed UC-MPC framework achieves a substantial reduction in online computational cost. In both vanilla and tube MPC, the optimization problem must be solved online at every sampling instant using the actual state of the system at the current time step as the initial condition or the constraint on the initial condition. In contrast, UC-MPC decouples the online optimization from the actual system evolution: the MPC problem is formulated solely based on the nominal dynamics in a closed-loop and can therefore be solved offline prior to control execution. This simplification is made possible by the integration of $\mathcal{L}_{1}$}AC, which ensures that the actual closed-loop system closely tracks the nominal system dynamics in (12).
@@
Overall, UC-MPC preserves the robustness guarantees of tube-based approaches while significantly reducing real-time computational demands by shifting the majority of the computation to the offline phase.

### v3

# Robust Adaptive MPC in the Presence of Nonlinear Time-Varying Uncertainties: An Uncertainty Compensation Approach

@@ Abstract
thanks: This work is supported by the Air Force Office of Scientific Research Grant (AFOSR) Grant AF FA9550-25-1-0274, the National Aeronautics and Space Administration (NASA) under Grant 80NSSC22M0070, and by the National Science Foundation (NSF) under Grants CMMI 2135925, CPS 2311085 and IIS 2331878. The material is this paper was partially presented at the 2024 American Control Conference (ACC), 10-12 July 2024, Toronto, ON, Canada. Corresponding author Pan Zhao.
@@ 2 Problem Statement
It is possible to extend the results to allow $w(t)$ to be state-dependent, $w(t,x)$, provided the boundedness assumption (3) holds. To simplify the exposition, we limit the derivations in this paper to time-dependence only.
@@ 3 Overview of the UC-MPC Framework
$\!\dot{x}(t)\!$ $=\!{A_{m}}x(t)\!+\!Bu_{\text{opt}}(t)\!+\!B(u_{\textup{a}}(t)\!+\!f(t,\!x(t)\!)\!)\!+\!B_{u}w(t),$
$y(t)$ $=Cx(t),\quad x(0)=x_{0},$ (8)
@@ 3.1 Equivalent Nominal System and Connection with Tube MPC
Why use two nominal systems? We first introduce the nominal system in its original form (9) because the performance analysis of $\mathcal{L}_{1}$}AC—specifically the derivation of its uniform performance bounds —relies on comparing this nominal system with the actual closed-loop dynamics, where their difference arises solely from system uncertainties. These performance bounds will be discussed in detail in Section 5. The transformation in (12) reveals a close connection between tube MPC and our proposed UC-MPC. In both frameworks, the feedback term $K_{x}(x-x_{\textup{n}})$ serves to keep the actual trajectory within a bounded neighborhood of the nominal one, while the adaptive component $u_{a}$ in UC-MPC actively compensates for matched uncertainties to further improve the performance. Owing to this adaptive compensation, the resulting deviation tube in UC-MPC is expected to be significantly tighter than that of the standard tube MPC. This reformulation also allows the MPC input $u_{\text{opt}}$ in (9) to be computed directly from (12), which defines a standard uncertainty-free MPC problem as in [22, 4, 6]. Consequently, the stability and recursive feasibility of the nominal MPC can be established using existing MPC analysis tools, which will be discussed in detail in Section 6.2. To summarize, the nominal system in (9) and control law in (7) serve as the foundation for all performance bounds analysis in Section 5, while the equivalent transformed form in (12) and control law in (13) are introduced for formulating the nominal MPC problem and highlighting the connection between UC-MPC and tube MPC [16], as well as for the stability and recursive feasibility analysis of the nominal MPC problem within the UC-MPC framework.
@@ 4 Unmatched Uncertainty Attenuation
Consider the following system:
@@
where $x(t)\in\mathbb{R}^{n}$ and $u(t)\in\mathbb{R}^{m}$ represent the state and input vectors, respectively, $w(t)\in\mathbb{R}^{p}$ represents the unmatched uncertainty, $z(t)$ denotes output channels whose bounds we aim to minimize, and matrices $A,B$, and $B_{u}$ are from (1). Note that when the matched uncertainty is perfectly eliminated, the original system (1) is identical to the system in (14a). With the feedback control law $u(t)=K_{x}x(t)$, we achieve the following closed-loop system:
@@
$\begin{bmatrix}(A\!+\!BK_{x})^{\top}P\!+\!P(A\!+\!BK_{x})\!+\!\lambda P&PB_{u}\\ B^{\top}_{u}P&-\mu I\end{bmatrix}<0,$ (16)
@@
Multiplying (18) by $\begin{bmatrix}\mathcal{V}^{-1}&0\\ 0&I\end{bmatrix}$ and its transpose from left and right gives (16), where $P=\mathcal{V}^{-1}$ and $K=\mathcal{Y}\mathcal{V}^{-1}$. Similarly, multiplying (19) by $\textup{diag}(\mathcal{V}^{-1},I,I)$ and its transpose from left and right gives (17). ∎
Thus, we can leverage Lemma 1 to achieve the internal stability and PPG bound $\beta$. Based on Lemma 2, we can derive matrices $\mathcal{Y}$ and $\mathcal{V}$ to minimize $\beta$ for a given $\lambda$ by solving the following LMI:
@@
By setting $K_{x}=\mathcal{Y}\mathcal{V}^{-1}$, we achieve the minimized PPG bound $\beta$ for system $H$ and less tightened constraint for channels in $z(t)$. Here, we have to pick a $\lambda$ first since 16 and 17 are only linear if we fix $\lambda$ [25]. For the selection of $\lambda$, a linear search between 0 to some predefined maximum value can be implemented for an improved performance.
@@
Lemma 2 provides one method to synthesize a state-feedback controller $u=K_{x}x$ to minimize the effect of bounded disturbances. There are other methods available in the literature, e.g., [1] and [9], both of which leverage inescapable ellipsoids to over-approximate the reachable sets of the disturbed system 14 and provide necessary and sufficient conditions for ensuring that system trajectories remain in prescribed ellipsoids. These methods can be seamlessly integrated into our proposed UC-MPC framework to replace the approach stated in Lemma 2, and can potentially lead to less conservative performance.
@@ 5 $\mathcal{L}_{1}$}AC with Uniform Performance Bounds
Given the uncertain system 8 subject to Assumption 1, the reference system 30 subject to the conditions 22b and 22a with a constant $\gamma_{1}>0$, and an $\mathcal{L}_{1}$}AC defined via 24, 25 and 26, subject to the sample time constraint 29, we have
@@
Given the reference system 30 and the nominal system 9, subject to Assumption 1, and the condition 22a, we have
@@ 5.1 Individual Performance Bounds
Consider the reference system 30 subject to Assumption 1, the nominal system 9, the transformed reference system 35 and transformed nominal system 39, obtained by applying 33 with any $\mathrm{T}_{x}^{i}$ satisfying 34. Suppose that 22a holds with some constants $\rho_{r}$ and $\left\lVert u_{\text{opt}}\right\rVert_{\mathcal{L}_{\infty}}$. Then, there exists a constant $\check{\rho}_{r}^{i}\leq\rho_{r}$ such that 41 holds with the same $\left\lVert u_{\text{opt}}\right\rVert_{\mathcal{L}_{\infty}}$. Furthermore, $\forall t\geq 0$,
@@
Consider the uncertain system 8 subject to Assumption 1, the nominal system 9, and the $\mathcal{L}_{1}$}AC defined via 24, 25 and 26, subject to the conditions 22b and 22a with constants $\rho_{r}$ and $\gamma_{1}>0$ and the sample time constraint 29. Suppose that for each $i\in\mathbb{Z}_{1}^{n}$, 41 holds with a constant $\check{\rho}_{r}^{i}$ for the transformed reference system 35 obtained by applying 33. Then, $\forall t\geq 0$, we have
@@
Lemma 6 introduces a technique to individually bound $x_{i}(t)-x_{\textup{n},i}(t)$ for each $i\in\mathbb{Z}_{1}^{n}$ and $u_{j}(t)-u_{\textup{n},j}(t)$ for each $j\in\mathbb{Z}_{1}^{m}$ through coordinate transformations. Furthermore, in the absence of unmatched uncertainty, by reducing $T$ and increasing the bandwidth of the filter ${\mathcal{C}}(s)$, the value of $\tilde{\rho}^{i}$ ($i\in\mathbb{Z}_{1}^{n}$) can be made arbitrarily small. This implies that the adaptive system states can be made arbitrarily close to the nominal system states. Additionally, the bounds on $u_{\textup{a},j}(t)$ and $u_{j}(t)-u_{\textup{n},j}(t)$ can be brought arbitrarily close to the bounds on the true matched uncertainty $f_{j}(t,x)$ for $x\in\mathcal{X}_{a}$ and each $j\in\mathbb{Z}_{1}^{m}$. We notice that the low-pass filter in $\mathcal{L}_{1}$}AC decides the tradeoff between performance and robustness. While reducing $T$ can lead to high-frequencies in the estimation loop, an appropriately selected bandwidth and order for the low-pass filter should prevent these high-frequencies to enter the system. Additional details on the role and systematic design of the filter can be found in [7].
@@ 5.2 Scaling for Improved Bounds
where $\bar{B}_{u}=B_{u}\Lambda$ and $\bar{w}$ = $\Lambda^{-1}w$ for some positive diagonal matrix $\Lambda$ s.t. $\left\lVert\bar{w}_{i}(t)\right\rVert_{\mathcal{L}_{\infty}}=b_{w}\;\forall i$. It follows that $B_{u}w=\bar{B}_{u}\bar{w}$, and thus the system (47) is identical to the original system (1), and the assumption in (4c) still applies to $\bar{w}(t)$. We can easily achieve the same results in Section 5.1 for the transformed system (47) but with smaller performance bounds. This is because during the calculation of the performance bounds in $\mathcal{L}_{1}$}AC, we only leverage $\mathcal{L}_{\infty}$ bound $b_{w}$. The additional information of individual bounds $b_{w_{j}}$ can help us design $\Lambda$, which makes the elements in $\bar{B}_{u}$ smaller than $B_{u}$ as $b_{w_{j}}\leq b_{w}$, and thus leading to smaller performance bounds for the transformed system (47) compared with the original system (1). In Section 7.1.5, we will show that with the transformation of $\bar{B}_{u}$ and $\bar{w}$ in (47), which we denote as scaling of $w(t)$, we achieve a smaller performance bound for every state and control input channel provided by $\mathcal{L}_{1}$}AC compared with the original system (1).
@@ 6 UC-MPC: Robust Adaptive MPC via Uncertainty Compensation
2:Calculate $\left\lVert u_{\text{opt}}\right\rVert_{\mathcal{L}_{\infty}}$ based on $u_{\textup{n}}(t)=K_{x}x_{\textup{n}}(t)+u_{\text{opt}}(t)$ with constraints $x_{\textup{n}}\in\mathcal{X}$ and $u_{\textup{n}}\in{\mathcal{U}}$ using Pontryagin set difference
@@
  • Design a low-pass filter in (21) that satisfies (22);
  • Apply the transformation 33 with $\mathrm{T}_{x}^{i}$ satisfying 34 to achieve $\check{\rho}_{r}^{i}$ that satisfies 41;
@@
The details are summarized in Algorithm 1. It is important to note that in step 3 of Algorithm 1 we also impose the constraint that both $x_{\textup{r}}(t)$ and $x(t)$ remain within $\mathcal{X}$ for all $t\geq 0$. These constraints can help reduce the size of the uncertainty that needs to be addressed and substantially decrease the conservatism of the proposed approach. Finally, we formulate an MPC problem for the nominal system 12 using tightened constraint sets.
@@ 6.1 Nominal MPC Formulation
where $T_{f}$ is the time horizon, $l_{f}$ denotes the custom terminal cost function, $l$ denotes the running cost function, and $\mathcal{X}_{f}$ is the custom terminal constraint set. Finally, we can apply the control $\bar{u}_{\text{opt}}$, defined as $\bar{u}_{\text{opt}}(\tau+t)=u^{*}(t)$, to the system, for $0\leq t\leq t_{\delta}$, where $u^{*}(\cdot)$ is the solution to the optimization problem 49 given $x_{\textup{n}}(\tau)$, and $t_{\delta}$ is a sufficiently small sampling time. It is worth noting that $\bar{u}_{\text{opt}}$ is derived solely from the nominal closed-loop dynamics, without relying on the actual system state affected by uncertainties.
@@ 6.4 Formal Guarantees of UC-MPC
Consider an uncertain system (1) subject to Assumption 1 with the state and input constraints in (6). Consider the control law (13), where $u_{a}$ is the control input of an adaptive controller designed following Algorithm 1, and $\bar{u}_{\text{opt}}$ is obtained from solving the nominal MPC problem 49. If Assumptions 2 and 3 hold, we have:
@@ 6.5 Computational Complexity Analysis
Compared with tube MPC formulations in [16, 10], the proposed UC-MPC framework significantly reduces the online computational burden. In tube MPC, the MPC optimization must be solved online at every sampling instant using the actual system state, and the nominal initial state is typically introduced as a decision variable constrained to be consistent with the actual state through the tube set. This increases the size and complexity of the optimization problem. In contrast, the MPC problem in (49) is solved online using only the nominal dynamics and nominal system state, without introducing the nominal initial state as a decision variable tied to the actual state. This simplification is enabled by the integration of $\mathcal{L}_{1}$}AC, which guarantees that the actual closed-loop system remains close to the nominal dynamics in (12). As a result, the optimization problem remains comparable in complexity to a vanilla MPC while yielding improved performance and robustness against model uncertainties.
@@
Overall, UC-MPC preserves the robustness guarantees of tube-based approaches while significantly reducing real-time computational effort.